OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -1 + 2 + 5 + 8 + ... + 44

OpenStudy (anonymous):

please help :/

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

well answer is 344. 2-(-1) = 3 5-2 =3 8-5 =3 the difference is 3 every time so just find every number after the 3rd number, then add them all together.